OpenStudy (anonymous):

What is the slope of the line that passes through (5,-3) and (-2,1)

OpenStudy (anonymous):

m= mx+b

OpenStudy (p0sitr0n):

slope m = (y2-y1)/(x2-x1) 1-(-3)/-2-5=4/-7

OpenStudy (anonymous):

\[slope=\frac{y_2-y_1}{x_2-x_1}\] Pute (x1, y1) =(5,-3) and (x2, y2) = (-2,1) into the formula.

OpenStudy (p0sitr0n):

no, m cant be equal to mx+b, m is the slope, not the line

OpenStudy (anonymous):

*put

OpenStudy (anonymous):

positron is right the slope is -4/7